# Enforcing asymptotic behavior with DNNs for approximation and regression in finance

### v1

@@ 1 Introduction
Often solutions to scientific problems and financial problems have known asymptotic behavior. In particular pricing functions in quantitative finance often have or are assumed to have linear asymptotes (Huge and Savine, 2020a, Section 2 in Appendix 4). If one represents pricing functions with a deep neural network (DNN) with activation functions with certain properties, these pricing functions will be asymptotically linear, but training the asymptotic part implicitly as part of the DNN presents challenges and requires special treatment (Huge and Savine, 2020a, Section 2 in Appendix 4). Often, unless specifically addressed in the training and setup, convergence of the DNN to the true asymptotic behavior can be slow or not even assured. Thus, one is interested in ways to use DNN to represent solutions while ensuring a given asymptotic form.
Antonov et al. (2020a, b) capture asymptotic behavior by a spline interpolator while in the interior, the solution is approximated by a linear combination of Gaussian kernels. The centers and widths of these kernels are given by artificial neural networks (Antonov et al. (2020a) calls this construction “Constrained Radial Layer” or “Gaussian Radial Layer”). They are constrained such that the resulting linear combination has zero asymptotes and is numerically zero in the asymptotic region beyond a certain distance.
@@ 2 Approach and Set-Up
with $s_{0}:=\frac{UI-LI}{UL-LL}$, $\tilde{LS}:=\frac{LS-s_{0}}{UL-LL}$, $\tilde{US}:=\frac{s_{0}-US}{UL-LL}$, and $a_{0}:=\frac{\tilde{US}-\tilde{LS}}{UL-LL}$. (Antonov et al. (2020a) present a more general version, but our form is enough for our purposes.) For alternative asymptotic forms, we would use these forms below $LL$ and above $UL$, and set $LS$ and $LI$ respective $US$ and $UI$ to the value and derivative of these forms at $LL$ and $UL$ and otherwise proceed in the same way. Antonov et al. (2020a) present an extension by multivariate splines for the multivariate box case and we could use their construction or alternative constructions for multivariate extensions of our approach as well.
Antonov et al. (2020a) ensure smooth pasting to the asymptotic forms by representing $nonasymptotic(x)$ as a linear combination of Gaussian kernels, with a more involved domain decomposition (asymptotic domain, "no-mans land", and inner domain). Here, we will use simpler forms.
@@
For testing examples where we know the function or the conditional expectation to be approximated, we can compute the error by comparing against that function and approximate the minimum loss by evaluating that loss function against generated samples. For the regression and conditional expectation setting, Cheridito and Gersey (2023a, b) offers ways to compute estimates and guarantees.
@@ 3 A first function approximation example
Now that we have defined the function, we generate input data for training using 50,000 randomly selected $x$ from a uniform distribution between -10 and 10 along with associated true function and derivative values $y$ and $\frac{dy}{dx}$. In a first approach, we learn the parameters in the asymptotic form through stochastic gradient descent or Adam Kingma and Ba (2014) at the same time the parameters of the neural network DNN are learned (“trainable parameters”). In a second approach, we use fixed given values for the parameters (“fixed parameters”). We tested using the true values of the asymptotic parameters (here, LS, LI, US, and UI) used to generate the samples as well as estimating the asymptotic parameters (LS, LI, US, and UI) through least squares in the asymptotic regions separately, possibly constraining the intercept to ensure good fitting of function values at LL and UL. We found that for our examples, the difference between using true values and using values determined with least square estimation was negligible. Thus, we present the results with parameters determined by least square estimation in the asymptotic regions for all our examples, since this would be also applicable if true parameters are not known.
@@ 6 Conclusion
While some approaches in the literature enforce asymptotic forms Antonov et al. (2020a, b), they represent functions in the non-asymptotic domain by a linear combination of Gaussian kernels (on top of the extended asymptotic form), requiring computation of many exponentials during training and evaluation. In this paper, we proposed and tested a simpler form where instead of Gaussian kernels, the product of a simple polynomial (which assures proper matching of asymptotic behavior) and an unconstrained DNN is used. This approach leads to a simple implementation and good results in our tests, on two example function approximation and one example regression problem. It thus seems to be a promising choice and candidate for implementation in cases where asymptotic behavior is known.
In future work, we plan to test further cases with known asymptotic behaviors, including the also encountered exponential forms. We intend to work on extensions to multiple dimensions. We plan to apply such asymptotic treatment in the regression setting in our PDML approach for parametric pricers and calibration Polala and Hientzsch (2023, 2024), and to the applications of the DML approach for regression to XVA Huge and Savine (2020a, b). We will also pursue the application of such constructions in further areas beyond the function approximation and regression setting, such as in stochastic control approaches Fathi and Hientzsch (2023); Hientzsch (2023), reinforcement learning Fathi and Hientzsch (2023); Hientzsch (2023), and deepBSDE approaches Hientzsch (2019, 2021); Yu et al. (2020, 2023).

### v2

@@ 1 Introduction
Often solutions to scientific problems and financial problems have known asymptotic behavior. In particular pricing functions in quantitative finance often have or are assumed to have linear asymptotes [HS20, Section 2 in Appendix 4]. If one represents pricing functions with a deep neural network (DNN) with activation functions with certain properties, these pricing functions will be asymptotically linear, but training the asymptotic part implicitly as part of the DNN presents challenges and requires special treatment [HS20, Section 2 in Appendix 4]. Often, unless specifically addressed in the training and setup, convergence of the DNN to the true asymptotic behavior can be slow or not even assured. Thus, one is interested in ways to use DNN to represent solutions while ensuring a given asymptotic form.
[AKP20, AKP20a] capture asymptotic behavior by a spline interpolator while in the interior, the solution is approximated by a linear combination of Gaussian kernels. The centers and widths of these kernels are given by artificial neural networks ([AKP20] calls this construction “Constrained Radial Layer” or “Gaussian Radial Layer”). They are constrained such that the resulting linear combination has zero asymptotes and is numerically zero in the asymptotic region beyond a certain distance.
@@ 2 Approach and Set-Up
with $s_{0}:=\frac{UI-LI}{UL-LL}$, $\tilde{LS}:=\frac{LS-s_{0}}{UL-LL}$, $\tilde{US}:=\frac{s_{0}-US}{UL-LL}$, and $a_{0}:=\frac{\tilde{US}-\tilde{LS}}{UL-LL}$. ([AKP20] present a more general version, but our form is enough for our purposes.) For alternative asymptotic forms, we would use these forms below $LL$ and above $UL$, and set $LS$ and $LI$ respective $US$ and $UI$ to the value and derivative of these forms at $LL$ and $UL$ and otherwise proceed in the same way. [AKP20] present an extension by multivariate splines for the multivariate box case and we could use their construction or alternative constructions for multivariate extensions of our approach as well.
[AKP20] ensure smooth pasting to the asymptotic forms by representing $nonasymptotic(x)$ as a linear combination of Gaussian kernels, with a more involved domain decomposition (asymptotic domain, "no-mans land", and inner domain). Here, we will use simpler forms.
@@
For testing examples where we know the function or the conditional expectation to be approximated, we can compute the error by comparing against that function and approximate the minimum loss by evaluating that loss function against generated samples. For the regression and conditional expectation setting, [CG21, CG23] offers ways to compute estimates and guarantees.
@@ 3 A first function approximation example
Now that we have defined the function, we generate input data for training using 50,000 randomly selected $x$ from a uniform distribution between -10 and 10 along with associated true function and derivative values $y$ and $\frac{dy}{dx}$. In a first approach, we learn the parameters in the asymptotic form through stochastic gradient descent or Adam [KB14] at the same time the parameters of the neural network DNN are learned (“trainable parameters”). In a second approach, we use fixed given values for the parameters (“fixed parameters”). We tested using the true values of the asymptotic parameters (here, LS, LI, US, and UI) used to generate the samples as well as estimating the asymptotic parameters (LS, LI, US, and UI) through least squares in the asymptotic regions separately, possibly constraining the intercept to ensure good fitting of function values at LL and UL. We found that for our examples, the difference between using true values and using values determined with least square estimation was negligible. Thus, we present the results with parameters determined by least square estimation in the asymptotic regions for all our examples, since this would be also applicable if true parameters are not known.
@@ 6 Conclusion
While some approaches in the literature enforce asymptotic forms [AKP20, AKP20a], they represent functions in the non-asymptotic domain by a linear combination of Gaussian kernels (on top of the extended asymptotic form), requiring computation of many exponentials during training and evaluation. In this paper, we proposed and tested a simpler form where instead of Gaussian kernels, the product of a simple polynomial (which assures proper matching of asymptotic behavior) and an unconstrained DNN is used. This approach leads to a simple implementation and good results in our tests, on two example function approximation and one example regression problem. It thus seems to be a promising choice and candidate for implementation in cases where asymptotic behavior is known.
In future work, we plan to test further cases with known asymptotic behaviors, including the also encountered exponential forms. We intend to work on extensions to multiple dimensions. We plan to apply such asymptotic treatment in the regression setting in our PDML approach for parametric pricers and calibration [PH23, PH24], and to the applications of the DML approach for regression to XVA [HS20, HS20a]. We will also pursue the application of such constructions in further areas beyond the function approximation and regression setting, such as in stochastic control approaches [FH23, H23], reinforcement learning [FH23, H23], and deepBSDE approaches [H19, H21, YHG20, YGH23].